# arXiv:2509.10756v2[quant-ph] 01 Mar 2026

### v2

@@ Abstract
We show that ensembles of deep neural networks, called deep ensembles, can be used to perform quantum parameter estimation while also providing a means for quantifying uncertainty in parameter estimates, which is a key advantage of using Bayesian inference for parameter estimation that is lost when using existing machine learning methods. We show that optimizing for both accurate parameter estimates and well calibrated uncertainty estimates does not lead to degradation in the former as opposed to only optimizing for accuracy. We also show that the drift detection capabilities of these ensemble models can be used to detect drift in the experimental data used during inference. These results suggest that such models could enable accurate, real-time parameter estimation with quantified uncertainty, making them promising candidates for deployment in experimental settings.
@@ I Introduction
The precise estimation of physical parameters is of central importance in quantum information science and physics more broadly [30, 52]. It is particularly relevant to the field of quantum metrology which itself has been instrumental for making discoveries in fundamental physics. For example squeezed-state based metrological methods have been used for gravitational wave detection [53, 50, 1] and dark matter axion searches [5, 7, 45]. Quantum parameter estimation procedures generally have three steps: i) a quantum probe system is prepared in some initial state $\rho_{0}$; ii) the unknown parameter(s) $\theta$ are encoded into the state of the probe system using some unitary transformation $U(\theta)\rho_{0}U^{\dagger}(\theta)$; iii) a measurement is performed on the probe system and the measurement result is used to extract information about the unknown parameter(s). This process is then repeated multiple times to gather more measurement data to improve the precision of the estimation [30].
@@
A popular approach for performing quantum parameter estimation is based on Bayesian inference [28, 16, 29]. This approach provides benefits such as uncertainty quantification for parameter estimates, a natural mechanism by which one can incorporate prior knowledge about the parameter into the analysis, and, perhaps most importantly, the ability to provide an optimal estimator of the parameters under consideration in the limit of a large enough number of measurements [16, 29]. However this approach also comes with significant drawbacks.
@@
Firstly Bayesian inference protocols generally require derivation of an analytic expression of the likelihood distribution that the measurement data is assumed to be generated from. This is not always feasible for many systems under ideal conditions and for simple systems this can become intractable when one attempts to model the effects of experimentally imperfections. This may be circumvented by use of so-called ‘likelihood-free’ methods such as approximate Bayesian computation (ABC) [46, 10] but this method presents it’s own challenges such as a drastic increase in sample complexity when the number of parameters being estimated increases, requiring access to a simulator which can generate samples from the likelihood distribution, and selection of a suitable discrepancy measure (and summary statistic) for the rejection sampling routine [36, 18, 46].
@@
Secondly Bayesian inference time can scale poorly as the dimensionality of the problem, i.e. number of parameters being estimated, increases. This is due to the random walk behavior used by common Markov chain Monte Carlo (MCMC) methods, such as Gibbs sampling, to explore the posterior distribution. This can lead to long convergence times when the the posterior distribution is high dimensional [18]. While more sophisticated Bayesian inference methods based on the Hamiltonian Monte Carlo algorithm are more robust to this they also introduce the additional overhead of requiring gradient computations to more effectively explore the posterior distribution [32, 18]. Bayesian inference time can also scale poorly when the number of measurement results used is large [17]. For these reasons real time estimation is made difficult.
@@ III.3 Hyperparameter Tuning
Ref. [43] trained single NNs on a dataset of $4\times 10^{6}$ trajectory-parameter pairs [44]. With the use of hyperparameter tuning we are able to achieve the same level of performance with only $1\%$ ($4\times 10^{4}$) of the original training data. This is also much lower than the amount of data needed for ABC which used 1000 trajectories generated for 100 different values of $\Delta$ which results in a total dataset size of $10^{5}$ [10, 12]. Note also that the ABC approach used longer trajectories with $N=200$ time delay measurements as opposed to the $N=48$ used in this work. We split the training data into training and validation sets with a $80\%$ and $20\%$ split respectively. A commonly used approach to train ensembles models is the use of bootstrap aggregation (bagging) which consists of training ensemble members on different bootstrapped samples of the original dataset. However, this is mainly advantageous for ensemble models comprised of weak learners, such as tree-based ensemble models, and worsens performance when used to training ensembles of strong learners like deep neural networks [33, 31]. We therefore did not utilize bagging in our experiments.
@@ III.5.2 Uncertainty quantification under data shift
IV Comparison to Approximate Bayesian Computation
@@ IV.1 Quantum system
where $\hat{a}$ and $\hat{b}$ are the annihilation operators of the optical and mechanical modes respectively, $\Omega$ is the laser Rabi frequency, $g$ is the optomechanical coupling strength, and $\omega_{M}$ is the mechanical frequency. The mechanical mode interacts with a thermal bath by both emitting and absorbing phonons while the optical mode leaks photons which are continuously detected. The master equation for this open system is given by
@@
IV.2 Predictive accuracy benchmark
@@
We train our deep ensemble to again minimize the Gaussian negative log likelihood loss. For the training data we use the data library from Ref. [10]. The data contains time delay trajectories of $80$ photon detections for 100 evenly spaced values of the detuning $\Delta$ across the range [10,0][-10,0] with $4,000$ trajectories for each unique $\Delta$. These data were generated using parameters We randomly split the library using an 80/20 split between training and testing data and stratify by $\Delta$ value to ensure each set is balanced and that we use a uniform prior for the ABC algorithm as done in [10]. Due to the nature of the dynamics of the system in the nonlinear regime, i.e. the system can generally not be assumed to reset to a consistent state with each photon emission like in the case of the TLS, the assumption of independence and identical distribution of the time delays in a given trajectory no longer holds. While these additional time delay correlations allow for potential improvement in the ability to perform sensing tasks [11] they also lead to the smoothed histogram neural network layer used previously no longer being optimal.
@@
Two tools that are more appropriate in this setting are recurrent neural network layers such as a long-short term memory layer or a one-dimensional convolutional layer. We use the latter as it is generally simpler to implement and more amenable to parallelization than recurrent architectures [51]. The specific architecture of the deep ensemble networks are shown in Table 1. We use an ensemble of $M=5$ networks with $\beta_{1}=0.87$, $\beta_{2}=0.93$, a learning rate of $10^{-3}$ for the Adam optimizer, and a batch size of $250$ over 50 epochs.
@@
Algorithm 1 Approximate Bayesian computation algorithm
2:Initialize empty histogram $H$ of $\theta$’s
[…]
4:while $i<\nu$ do
6:  if $\delta(S((D^{\prime},\theta)),S(D))\leq\epsilon$ then
8:  end if
9:  $i\leftarrow i+1$
10:end while
11:Normalize $H$
We draw $\nu=10^{5}$ samples and $\epsilon=9$ as this was found to be more than enough for the method to converge for estimation of the same parameter and system in [10]. We also use the same histogram based summary statistic and $L_{2}$ discrepancy measure from [10] and similarly take the mean of the ABC posterior as our estimator of the parameter.
Fig. 8 shows the performance of the two methods on our test set. We can see that for most values of $\Delta$ the deep ensemble achieves lower error than ABC and achieves competitive accuracy in all other cases. Interestingly, for the $\Delta_{n=3}$ regime we see that the ensemble greatly outperforms the ABC method suggesting it makes better use of the information encoded in the three-photon correlations present in the trajectory data.
V Computational Efficiency
Another advantage provided by NNs is their computational efficiency in terms of inference time and memory overhead particularly when compressed for deployment on edge devices such as FPGAs. The deep ensembles discussed thus far have a memory size of roughly 1.1 megabtyes. This can be further compressed via the use of model quantization which converts the 32-bit floating point weights of the model to 8-bit integers. This reduces the size of our models to roughly 8.8 kilobytes. This is valuable in cases where one wishes to deploy these models to edge devices with limited memory to perform real time inference and does not lead to any noticeable deterioration of model performance as can be seen in Fig. 9(a). Other memory optimization methods, such as weight pruning, can be performed if further memory efficiency is required.
We also compare inference times by measuring the wall clock time of a trained quantized deep ensemble, the ABC method and an optimized likelihood-based Hamiltonian Monte Carlo (HMC) Bayesian inference routine for estimating $\Delta$ given delay measurements from different numbers of trajectories. We use the state-of-the-art No U-Turn Sampler (NUTS) provided by the PyMC framework as our specific HMC method [24, 2]. For inference with NUTS we use 4 chains and draw 1,000 posterior samples per chain which we parallelize over 4 cores. For inference with ABC we use the same implementation of the algorithm used in the previous section which leverages NumPy array vectorization when possible to ensure fair comparison. The average inference time for the three methods over 20 different runs of the inference on an AMD Ryzen 9 3900X 12-Core Processor are shown in Fig. 9(b). We can see that the highly parallelizable nature of neural network inference, i.e. batched inference, allows for orders of magnitude faster inference speed over the NUTS and ABC method despite the need to perform a forward pass for each of the 10 NNs in the deep ensemble. The ability to parallelize inference with a neural network, due to the fact that the network is carrying out simple linear algebraic operations, contrasts with the inherently sequential nature of MCMC based Bayesian inference due to the construction of the Markov chains.
VI Conclusion
We have proposed a method for using NNs to perform parameter estimation which provides uncertainty estimates in addition to point estimates of the parameters. We showed that these models provide good parameter estimates despite being optimized for both predictive accuracy and well calibrated uncertainty estimation. We also demonstrated our scheme is applicable to more complex systems in which the measurement results are not independent and identically distributed, unlike the case of the TLS studied in much of the present work, where recurrent or convolutional NNs become more appropriate to use over the permutation-invariant architectures discussed initially. Furthermore, our method can, in principle, be generalized to $n$-dimensional estimation problems by only adding an additional $O(n)$ neurons in the output layer of the network. We leave this for future work along with the exploration of other NN architectures, such as deep sets, that possess permutation invariance and the ability to naturally handle inputs of varying length [55]. In this work we performed hyperparameter tuning on a single model and used the optimal hyperparameters from that search to train the deep ensemble due to time constraints. An interesting question to answer would be if hyperparameter tuning on the ensemble level, while being more computationally expensive, could deliver even better results. Finally, in this work we considered ensembles in which all models have the same architecture but this is by no means a necessity. Further exploration of how using ensembles of networks with different architectures, e.g. different activation functions and numbers of hidden layers, would impact ensemble performance on the parameter estimation task would be useful, in particular with respect to how this affects the uncertainty estimates as this approach may better capture epistemic (model) uncertainty.
Acknowledgements.
Appendix A Bayesian Inference
$P(\theta|D)=\frac{P(D|\theta)P(\theta)}{P(D)}$ (13)
where $P(D|\theta)$ is the likelihood, $P(\theta)$ is the prior distribution over parameter values which can be chosen to incorporate our prior knowledge about the parameter, and $P(D)$ is called the evidence or the marginal likelihood [32, 18]. The Bayesian estimator, $\hat{\theta}$, of $\theta$ is chosen to be either the mean, median or mode of the posterior with the posterior mean being the most common choice [14]. In our work we use the posterior mean $\mathbb{E}[P(\theta|D)]$. The delay time probability distribution, $w(\tau;\Delta,\Omega,\gamma)$, which gives the probability density for a particular delay time $\tau$ is given by [43]:
$P(D|\theta)=\prod_{i=1}^{N}w(\tau_{i};\Delta,\Omega,\gamma).$ (15)
For our prior we use a uniform distribution with support on the range $0,2.1\gamma]$.
Appendix B Hyperparameter Tuning Details
Hyperparameter Search Space
Number of histogram bins $(200,710)$
Loss function {RMSE, MSLE}
Learning rate $(10^{-5},5\times 10^{-3})$
$\beta_{1}$ of Adam optimizer $(0.8,0.999)$
$\beta_{2}$ of Adam optimizer $(0.8,0.999)$
Number of epochs $(50,500)$
Batch size $\{2^{6},2^{7},2^{8},2^{9},2^{10},2^{11}\}$
Dropout probability $(0,0.2)$
Early stopping patience $(4,10)$

### v3

# Quantum parameter estimation with uncertainty quantification from continuous measurement data using neural network ensembles

@@ Abstract
We show that ensembles of deep neural networks, called deep ensembles, can be used to perform quantum parameter estimation while also providing a means for quantifying uncertainty in parameter estimates, which is a key advantage of using Bayesian inference for parameter estimation that is lost when using existing machine learning methods. We show that optimizing for both accurate parameter estimates and well calibrated uncertainty estimates does not lead to degradation in the former as opposed to only optimizing for accuracy. We also show that the drift detection capabilities of these ensemble models can be used to detect drift in the experimental data used during inference. This approach is also shown to provide much faster inference time than both likelihood-based and likelihood-free Bayesian inference. These results suggest that such models could enable accurate, real-time parameter estimation with quantified uncertainty, making them promising candidates for deployment in experimental settings.
@@ I Introduction
The precise estimation of physical parameters is of central importance in quantum information science and physics more broadly [32, 54]. It is particularly relevant to the field of quantum metrology which itself has been instrumental for making discoveries in fundamental physics. For example, squeezed-state based metrological methods have been used for gravitational wave detection [55, 52, 1] and dark matter axion searches [5, 7, 47]. Quantum parameter estimation procedures generally have three steps: i) a quantum probe system is prepared in some initial state $\rho_{0}$; ii) the unknown parameter(s) $\theta$ are encoded into the state of the probe system using some unitary transformation $U(\theta)\rho_{0}U^{\dagger}(\theta)$; iii) a measurement is performed on the probe system and the measurement result is used to extract information about the unknown parameter(s). This process is then repeated multiple times to gather more measurement data to improve the precision of the estimation [32].
@@
A popular approach for performing quantum parameter estimation is based on Bayesian inference [30, 17, 31]. This approach provides benefits such as uncertainty quantification for parameter estimates, a natural mechanism by which one can incorporate prior knowledge about the parameter into the analysis, and, perhaps most importantly, the ability to provide an optimal estimator of the parameters under consideration in the limit of a large enough number of measurements [17, 31]. However, this approach also comes with significant drawbacks.
@@
Firstly, Bayesian inference protocols generally require derivation of an analytic expression of the likelihood distribution that the measurement data is assumed to be generated from. This is not always feasible for many systems under ideal conditions and for simple systems this can become intractable when one attempts to model the effects of experimentally imperfections. This may be circumvented by use of so-called ‘likelihood-free’ methods such as approximate Bayesian computation (ABC) [48, 11] but this method presents it’s own challenges such as a drastic increase in sample complexity when the number of parameters being estimated increases, requiring access to a simulator which can generate samples from the likelihood distribution, and selection of a suitable discrepancy measure (and summary statistic) for the rejection sampling routine [38, 19, 48].
@@
Secondly, Bayesian inference time can scale poorly as the dimensionality of the problem, i.e. number of parameters being estimated, increases. This is due to the random walk behavior used by common Markov chain Monte Carlo (MCMC) methods, such as Gibbs sampling, to explore the posterior distribution. This can lead to long convergence times when the the posterior distribution is high dimensional [19]. While more sophisticated Bayesian inference methods based on the Hamiltonian Monte Carlo algorithm are more robust to this they also introduce the additional overhead of requiring gradient computations to more effectively explore the posterior distribution [34, 19]. Bayesian inference time can also scale poorly when the number of measurement results used is large [18]. For these reasons real time estimation is made difficult.
@@ III.3 Hyperparameter Tuning
Ref. [45] trained single NNs on a dataset of $4\times 10^{6}$ trajectory-parameter pairs [46]. With the use of hyperparameter tuning we are able to achieve the same level of performance with only $1\%$ ($4\times 10^{4}$) of the original training data. This is also much lower than the amount of data needed for ABC which used 1000 trajectories generated for 100 different values of $\Delta$ which results in a total dataset size of $0^{5} [11$, 13]. Note also that the ABC approach used longer trajectories with $N=200$ time delay measurements as opposed to the $N=48$ used in this work. We split the training data into training and validation sets with a $80\%$ and $20\%$ split respectively. A commonly used approach to train ensembles models is the use of bootstrap aggregation (bagging) which consists of training ensemble members on different bootstrapped samples of the original dataset. However, this is mainly advantageous for ensemble models comprised of weak learners, such as tree-based ensemble models, and worsens performance when used to training ensembles of strong learners like deep neural networks [35, 33]. We therefore did not utilize bagging in our experiments.
@@ III.5.2 Uncertainty quantification under data shift
III.6 Posterior approximation
In addition to providing uncertainty estimates another benefit provided by our ensemble approach, which is similarly unavailable for currently existing single network approaches, is the ability to provide a smooth probability distribution over possible parameter values similar to the Bayesian inference approach. Naturally one may ask, given the uniform prior used in the Bayesian analysis and the universal approximation property of Gaussian mixtures, if we can expect the ensembles output distribution to have good fidelity with the Bayesian posterior. We provide an example of the Bayesian posterior and the ensemble mixture distribution for a single trajectory in Fig. 8 which suggests that this is indeed possible.
Figure 8: Posterior distribution computed with a uniform prior over $\Delta\in[0,5\gamma]$ and mixture distribution learned by the ensemble for an example trajectory generated with $\Delta\approx 1$ and $\Omega=1$.
To examine this further we compute the average fidelity between the Bayesian posterior distribution and the distribution learned by the ensemble over 100 random trajectories for each unique value of $\Delta$ in our test set. As done in Ref. [11] we compute the fidelity between the analytic posterior and the mixture distribution learned by the ensemble as the Bhattacharyya coefficient
$F(p_{A}(\theta),p_{B}(\theta))=\int_{\Theta}\sqrt{p_{A}(\theta)p_{B}(\theta)}d\theta.$ (9)
Figure 9: Average fidelity between Bayesian posterior and ensemble Gaussian mixture distributions over 100 random trajectories per $\Delta$ value.
The results are shown in Fig. 9. We can see that for the majority of $\Delta$ values the distribution learned by the ensemble has strong overlap with the Bayesian posterior. The fidelity degrades slightly near resonance values $\Delta\approx 0$ as the uniform prior leads to an abrupt vanishing of the posterior density for values of $\Delta<0$ leading to a non-smoothly varying distribution near these values unlike the one always produced by the ensemble. Additionally, these values being close to the limit of the range of $\Delta$ values, $0,5\gamma]$, used in the training data of the ensemble likely exacerbates the discrepancies.
III.7 Multi-parameter estimation
Here we demonstrate our approach applied to the simultaneous estimation of the Rabi frequency $\Omega$ and detuning $\Delta$. We train an ensemble of $M=10$ networks on $16\times 10^{4}$ trajectories of length $N=48$ which were generated using random values of $\Delta\in[0,3\gamma]$ and $\Omega\in[0.25\gamma,5\gamma]$ drawn uniformly from their respective intervals. We compare our ensembles performance to that of standard Bayesian inference performed using nested sampling [8, 16]. To avoid the prohibitively long inference time on our computational resources associated with the two-dimensional Bayesian inference routine we evaluate the trained deep ensemble on the test set of trajectories from Ref. [45] whose precomputed Bayesian point estimates for the test set trajectories are made available at [46]. The test set contains $16\times 10^{6}$ trajectories to ensure convergence of the computed RMSEs and a uniform prior was used over the interval from which the training set was generated from to ensure fair comparison. Fig. 10 shows the RMSE achieved by both the deep ensemble and Bayesian inference. We can see that both approaches perform comparably with the deep ensemble often achieving a lower RMSE particularly for parameter values close to the boundaries of the parameter space under consideration. This shows that not only can the deep ensemble exceed the inference speed of likelihood-based Bayesian inference but that it can also achieve better performance in terms of estimation accuracy.
Figure 10: Contour plots of average RMSE achieved by the deep ensemble and Bayesian inference via nested sampling on test set.
III.8 Comparison to Approximate Bayesian Computation
@@ III.8.1 Quantum system
where $\hat{a}$ and $\hat{b}$ are the annihilation operators of the optical and mechanical modes respectively, $\Omega$ is the laser Rabi frequency, $g$ is the optomechanical coupling strength, and $\omega_{M}$ is the mechanical frequency. The mechanical mode interacts with a thermal bath by both emitting and absorbing phonons while the optical mode leaks photons which are continuously detected. The master equation for the open system is given by
@@
III.8.2 Predictive accuracy benchmark
@@
We train our deep ensemble to again minimize the Gaussian negative log likelihood loss. For the training data we use the data library from Ref. [11]. The data contains time delay trajectories of $80$ photon detections for 100 evenly spaced values of the detuning $\Delta$ across the range [10,0][-10,0] with $4,000$ trajectories for each unique $\Delta$. We randomly split the library using an 80/20 split between training and testing data and stratify by $\Delta$ value to ensure each set is balanced and that we use a uniform prior for the ABC algorithm as done in [11]. Due to the nature of the dynamics of the system in the nonlinear regime, i.e. the system can generally not be approximated as a coherent state and assumed to reset to a consistent state with each photon detection like in the case of the TLS, the assumption of independence and identical distribution of the time delays in a given trajectory no longer holds. While these additional time delay correlations allow for potential improvement in the ability to perform sensing tasks [12] they also lead to the smoothed histogram neural network layer used previously no longer being optimal.
@@
Two tools that are more appropriate in this setting are recurrent neural network layers such as a long-short term memory layer or a one-dimensional convolutional layer. We use the latter as it is generally simpler to implement and more amenable to parallelization than recurrent architectures [53]. The specific architecture of the deep ensemble networks are given in Appendix C. We use an ensemble of $M=10$ networks with $\beta_{1}=0.87$, $\beta_{2}=0.93$, a learning rate of $10^{-3}$ for the Adam optimizer, and a batch size of $250$ over 50 epochs.
Algorithm 1 Approximate Bayesian computation algorithm
2:Initialize empty histogram $H$ of $\theta$’s
[…]
4:while $i<\nu$ do
6:  if $\delta(S((D^{\prime},\theta)),S(D))\leq\epsilon$ then
8:  end if
9:  $i\leftarrow i+1$
10:end while
11:Normalize $H$
We draw $\nu=10^{5}$ samples and set $\epsilon=9$ as this was found to be more than enough for the method to converge for estimation of the same parameter and data in [11]. We also use the same histogram based summary statistic and $L_{2}$ discrepancy measure from [11] and similarly take the mean of the ABC posterior as our estimator of the parameter.
Fig. 11 shows the performance of the two methods on our test set. We can see that for most values of $\Delta$ the deep ensemble achieves lower error than ABC and achieves competitive accuracy in all other cases. Interestingly, for the $\Delta_{n=3}$ regime we see that the ensemble greatly outperforms the ABC method suggesting it makes better use of the information encoded in the three-photon correlations present in the trajectory data. Note also that the inference time for the the ABC method is also about two orders of magnitude larger as we shall see in the next section.
IV Computational Efficiency
Here we more closely examine the advantage provided by NNs in terms of their computational efficiency with respect to inference time and memory overhead particularly when compressed for deployment on edge devices such as FPGAs. The deep ensembles discussed thus far have a memory size of roughly 1.1 megabtyes. This can be further compressed via the use of model quantization which converts the 32-bit floating point weights of the model to 8-bit integers. This reduces the size of our models to roughly 8.8 kilobytes. This contrasts with a method such as ABC which requires the storage of the data library during inference time and thus has a space complexity which is linear in the size of the dataset. This is valuable in cases where one wishes to deploy these models to edge devices with limited memory to perform real time inference and does not lead to any noticeable deterioration of model performance as can be seen in Fig. 12(a). Other memory optimization methods, such as weight pruning, can be performed if further memory efficiency is required.
We also compare inference times by measuring the wall clock time of a trained quantized deep ensemble, the ABC method and an optimized likelihood-based Hamiltonian Monte Carlo (HMC) Bayesian inference routine for estimating $\Delta$ given delay measurements from different numbers of trajectories. We use the state-of-the-art No U-Turn Sampler (NUTS) provided by the PyMC framework as our specific HMC method [26, 2]. For inference with NUTS we use 4 chains and draw 1,000 posterior samples per chain which we parallelize over 4 cores. For inference with ABC we use the same implementation of the algorithm used in the previous section which leverages NumPy array vectorization [23] when possible to ensure fair comparison. The average inference time for the three methods over 20 different runs of the inference on an AMD Ryzen 9 3900X 12-Core Processor are shown in Fig. 12(b). We can see that the highly parallelizable nature of neural network inference, i.e. batched inference, allows for orders of magnitude faster inference speed over the NUTS and ABC method despite the need to perform a forward pass for each of the 10 NNs in the deep ensemble. The ability to parallelize inference with a neural network, due to the fact that the network is carrying out simple linear algebraic operations, contrasts with the inherently sequential nature of MCMC based Bayesian inference, due to the construction of the Markov chains, and ABC based Bayesian inference due to the rejection sampling routine. This shows that deep ensembles are better equipped for real time quantum parameter estimation tasks while still retaining the uncertainty quantification capabilities of Bayesian inference.
V Conclusion
We have proposed a method for using NNs to perform parameter estimation which provides uncertainty estimates in addition to point estimates of the parameters. We showed that these models provide good parameter estimates despite being optimized for both predictive accuracy and uncertainty estimation. We also show that they provide well calibrated uncertainty estimates and good approximations of the Bayesian posterior distribution. We also demonstrated our scheme is applicable to more complex systems in which the measurement results are not independent and identically distributed, unlike the case of the TLS studied in much of the present work, where recurrent or convolutional NNs become more appropriate to use over the permutation-invariant architectures discussed initially. Furthermore, our method can, in principle, be generalized to $n$-dimensional estimation problems by only adding an additional $O(n)$ neurons in the output layer of the network as opposed to classification based approaches [40]. We have demonstrated this for the $n=2$ case in the present work. We leave higher dimensional extensions for future work along with the exploration of other NN architectures, such as deep sets, that possess permutation invariance and the ability to naturally handle inputs of varying length [57].
In this work we performed hyperparameter tuning on a single model and used the optimal hyperparameters from that search to train the deep ensemble due to time constraints. An interesting question to answer would be if hyperparameter tuning on the ensemble level, while being more computationally expensive, could deliver even better results. We also only considered ensembles in which all models have the same architecture but this is by no means a necessity. Further exploration of how using ensembles of networks with different architectures, e.g. different activation functions and numbers of hidden layers, would impact ensemble performance on the parameter estimation task would be useful, in particular with respect to how this affects the uncertainty estimates as this approach may better capture epistemic (model) uncertainty. Finally, since photon counting is not necessarily the most optimal measurement in terms of the quantum Fisher information (QFI) [45] it would be interesting to explore how the performance and data efficiency of these models changes when the QFI of the continuous measurement is optimized [9].
Acknowledgements.
Appendix A Bayesian Inference
$P(\theta|D)=\frac{P(D|\theta)P(\theta)}{P(D)}$ (14)
where $P(D|\theta)$ is the likelihood, $P(\theta)$ is the prior distribution over parameter values which can be chosen to incorporate our prior knowledge about the parameter, and $P(D)$ is called the evidence or the marginal likelihood [34, 19]. The Bayesian estimator, $\hat{\theta}$, of $\theta$ is chosen to be either the mean, median or mode of the posterior with the posterior mean being the most common choice [15]. In our work we use the posterior mean $\mathbb{E}[P(\theta|D)]$. The delay time probability distribution for the TLS, $w(\tau;\Delta,\Omega,\gamma)$, which gives the probability density for a particular delay time $\tau$ is given by [45]:
$P(D|\theta)=\prod_{i=1}^{N}w(\tau_{i};\Delta,\Omega,\gamma).$ (16)
For our prior we use a uniform distribution with support on the range the training data was genereated from for all experiments with the TLS.
Appendix B Hyperparameter Tuning Details
Hyperparameter Search Space
Number of histogram bins $(200,710)$
Loss function {RMSE, MSLE}
Learning rate $(10^{-5},5\times 10^{-3})$
$\beta_{1}$ of Adam optimizer $(0.8,0.999)$
$\beta_{2}$ of Adam optimizer $(0.8,0.999)$
Number of epochs $(50,500)$
Batch size $\{2^{6},2^{7},2^{8},2^{9},2^{10},2^{11}\}$
Dropout probability $(0,0.2)$
Early stopping patience $(4,10)$
Appendix C Convolutional neural network architecture
Here we provide the specific architecture of the convolutional neural network used in Section III.8 in Table 2.